# Implementing Fairness in AI Classification: The Role of Explainability

### v2

@@ Abstract
In this paper, we propose a philosophical and experimental investigation of the problem of fairness in AI classification. We argue that implementing fairness in AI classification involves more work than just operationalizing a fairness metric. It requires establishing the explainability of the classification model chosen and of the principles behind it. Specifically, it involves making the training processes transparent, determining what outcomes the fairness criteria actually produce, and assessing their trade-offs by comparison with closely related models that would lead to a different outcome. To exemplify this methodology, we trained a model and developed a tool for disparity detection and fairness interventions, the package FairDream. While FairDream is set to enforce Demographic Parity, experiments reveal that it fulfills the constraint of Equalized Odds. The algorithm is thus more conservative than the user might expect. To justify this outcome, we first clarify the relation between Demographic Parity and Equalized Odds as fairness criteria. We then explain FairDream’s reweighting method and justify the trade-offs reached by FairDream by a benchmark comparison with closely related GridSearch models. We draw conclusions regarding the way in which these explanatory steps can make an AI model trustworthy.
Keywords: Explainable AI ; Fairness ; Equalized Odds ; Demographic Parity; Calibration; GridSearch ; Trust
@@ 1 Introduction
Determining whether an algorithm provides ‘fair’ predictions is a challenging task, although one that is of central practical and ethical importance for society. The difficulties pertain in part to the fact that different notions of fairness exist and compete with each other. On the theoretical side, plethora of statistical metrics exist to quantify the extent to which the predictions of an AI model meet the user’s expectation on fairness, in particular regarding a minority that must not be harmed, [Hellman, 2020] [Mehrabi et al., 2021] [Wan et al., 2023]. Yet, attempts to compare these algorithmic fairness metrics face impossibility theorems, implying that the various fairness criteria proposed cannot be jointly satisfied in all cases (viz. [Kleinberg et al., 2016, Hellman, 2020, Hedden, 2021]).
@@
In this paper, we propose to look at the problem of defining a transparent and explainable AI system for inequality detection and correction. As with other approaches to AI fairness developed in diverse domains, including medicine [Chen et al., 2023], law [Lagioia et al., 2023], or mortgage lending [Lee and Floridi, 2021], we anchor our philosophical analysis to a particular case study, to draw general lessons from the way in which bias can be detected and corrected using AI. Specifically, we trained an AI model and developed our own fairness package (FairDream) to detect inequalities and then to correct for them. As experiments with FairDream reveal, however, the settings proposed by the system don’t necessarily produce expected outcomes, and so they call for an explanation in order to make the system not just usable, but also transparent, and trustworthy.
The way FairDream works is by implementing certain criteria for fairness correction. In relation to that, we draw attention to a tension between two fairness criteria: one promoting the same overall positive rate across groups (Demographic Parity), another recommending to equalize true positive rates and false positive rates across groups (Equalized Odds). In the literature on algorithmic fairness, Equalized Odds is also opposed to the notion of Calibration ([Mayson, 2018, Long, 2021, Hedden, 2021, Barocas et al., 2023]), often leaving Demographic Parity as a distant contender. However, our reasons to focus on the opposition between Equalized Odds and Demographic Parity rather than with Calibration are twofold.
@@
First of all, demands of Demographic Parity are real in various segments of society, in relation to gender, age, or ethnicity. Demographic Parity has thus been interpreted as a legal requirement, e.g. on EU Non-Discrimination Law [Wachter et al., 2020] or for the risk assessment mandated by Pennsylvania [Selbst et al., 2019]. Secondly, while we trained an AI model to enforce Demographic Parity, it eventually converged on the enforcement of Equalized Odds instead, in ways that are not just coincidental. Our experiments reveal that it is a property of FairDream to fulfill fairness objectives which are conditional on the ground truth (Equalized Odds), even when the user wants to achieve unconditional equality (Demographic Parity). While this may be seen as an anomaly, we first explain this property of FairDream, and then we use it to propose an argument for fairness metrics conditioned on true labels. FairDream is intended to enable a user to perform fairness interventions. We explain the way in which the criterion of Equalized Odds constrains such interventions, in particular by comparing it to similar in-processing methods that forego this constraint. From this comparison, we draw the lesson that one way of making an AI system trustworthy is to provide a comparison and contrastive explanation of its results with minimally modified versions of the system that produce significantly different outcomes.
@@
In Sections 2 and 3, we first present our training data and the package FairDream. Section 4 introduces the specific property of FairDream of enforcing Equalized Odds, and Section 5 proposes to explain this property by comparison with a closely related approach, the so-called GridSearch method, which can diverge from true labels in order to enforce Demographic Parity. In Section 6, finally, we present some arguments in favor of a metric conditioned on true labels. Firstly, we discuss how FairDream performs in terms of calibration, and we explain calibration is not a relevant criterion from our perspective, including when it is interpreted as a requirement of equal precision relative to a threshold. Then, we respond to the suspicion that the criterion used might be too conservative. We argue that a more revisionary notion of fairness like Demographic Parity too can be enforced, but provided its effects are carefully controlled for. Finally, in Section 7 we conclude with broader lessons regarding the way in which the explanatory steps taken here make the model trustworthy.
@@ 2 Age disparities in the evaluation of income
For our experiment, we thus manipulated the well-known Census dataset, which aggregates data from 48,842 US citizens in 1994.111https://www.kaggle.com/datasets/uciml/adult-census-income. The dataset includes 14 characteristics such as age, work class, years of education, marital status, and geographical origins. A fifteenth column, finally, displays a binary encoding of whether the agent earns less or more than \$50,000 a year (Cf. Figure 1).
@@
Using the Census dataset, we thus built a machine-learning model whose purpose is to minimize the error between its predictions and the actual income of people. Because the practitioner can choose different learners and different metrics of statistical performance, we built a tree-based AI model with the XGBoost library and evaluated it using the standard ROC-AUC measure. The models of trees ensembles as their evaluation through ROC-AUC are, indeed, broadly used in binary classification on tabular data [Shwartz-Ziv and Armon, 2022]. [^3] An XGBoost classifier, – for eXtreme Gradient Boosted trees,– is an ensemble of trees. Each tree tries to find the best features (splitting populations on their values) to distinguish between individuals earning more or less than \$50,000. Boosting relies on a sequence of weak trees, progressively correcting each others to reach the best predictor. In XGBoost, the computation is made faster through gradient approximation of the errors of the trees, and techniques such as tree-pruning are used to find new trees faster [Chen and Guestrin, 2016]. For stability of the following experiments, we reduced the architecture of the gradient boosted trees to a minimal tuning. By default, we set the model to 1,000 estimators, the maximal depth of each tree being 3 splits on features.
@@ 3 The package FairDream
How can we assess whether an AI model in charge of predicting level of income is fair before deploying it? To answer this question, we implemented our own algorithmic fairness package (FairDream). In FairDream, the main purpose is to give lay users, without any background in data-science or in theories of fairness, an insight into the baseline AI model, and subsequently, a handle on the process of correction.
@@
FairDream takes as input an already trained predictor (the baseline XGBoost we trained) and an indication of the users’ preference for fairness regarding the classificatory variable of interest (i.e. how much the user wants to impose parity regarding the percentage of individuals labeled as “\$$\geq\$50k$” by the XGBoost across groups). The procedure implemented in FairDream involves two steps, a step of detection and a step of correction.
@@
On the sample of Census used for training, FairDream  warns for a ratio of 1 to 5 in the model’s predictions: for example, for the 17-29 year-old individuals, the portion of the sub-population selected is 12%, against 66% for 29-37 year-old (Cf. Table 1), thereby triggering an alert.
@@
In our experiment, the stat_score is the indicator of predictive performance chosen by the user, such as ROC-AUC. The fair_score is maximized when the indicators of fairness determined by the user are equalized across groups, such as the overall positive rates for Demographic Parity. We implemented it as a weighted linear sum of the differences of fairness indicators between groups. The best model is then selected as the one with the highest ${\tt trade\_off\_score}=1/3*{\tt stat\_score}+2/3*{\tt fair\_score}$, if for instance the user prefers fairness over accuracy.
@@
Table 1 shows the predictions of the baseline model at the first step of the procedure, while Table 2 shows the predictions of the model selected as best by FairDream at the second step. Each table highlights the rates of positive predictions (overall positive, true positive, false positive) of the corresponding model’s predictions for two ages groups. The model is evaluated on each age group, first at the global level, then looking at the individuals having the true label ‘\$$\geq\$50,000$’, finally looking at the individuals having the true label ‘\$$\geq\$50,000$’.
At the global level, comparing Table 2 and Table 1 shows that the disparity between age groups in the corrected model is indeed less than 1 to 5, but the correction seems surprisingly small (16 % vs 59 % instead of 12 % vs 66 %), leaving a gap whose ratio is still of 1 to 3.7. When comparing by true labels, however, we do see an important correction, since true positive rates and false positive rates come out nearly equal (88% to 89%, 3% to 4%), unlike in the baseline model (68% to 90%, and 2% to 20%). This suggests that the model is not fulfilling Demographic Parity, but a different criterion, known as Equalized Odds. We review the difference between them in the next section.
@@ 4 Demographic Parity vs Equalized Odds
Adopting the taxonomy of [Wachter et al., 2020], we may first interpret the desire to diminish the gap between older and younger clients as an unconditional wish, namely to predict equal proportions of 17-29 and 29-37 year-old as earning over \$50,000, regardless of their true level of income. That would be translated into an objective of “Demographic Parity” [Dwork et al., 2012]: [^7] We notate $Y$ whether an individual actually earns over \$50k, and $\hat{Y}$ the prediction of the model; $A$ and $B$ respectively represent whether an individual is in the 17-29-year old or in the 29-37 year-old category.
@@
This is unexpected, since in this experiment we actually asked for the overall percentages of clients earning over \$50,000 to be predicted as equal across ages by the FairDream model. Mathematically, this objective corresponds to Demographic Parity. Yet, we see that Demographic Parity is not achieved by the FairDream  model. Instead, it actually produces a different result, namely Equalized Odds. Is it an accident? Or is it rather an emergent feature of the model, which FairDream can help us to rationalize? We proceed to clarify this issue in the next section.
@@ 5.1 An in-processing method
Once we set the percentage of clients predicted as earning over \$50k to be equal across ages (Demographic Parity), the fairness techniques can equalize them before, during, or after training of the model [Alves et al., 2021]. Although they are implemented in diverse ways, these techniques mainly rely on the following architectural principles:
@@ 5.2 Reweighting for transparent processing
In our attempt to make the method as transparent as possible to lay users, our method uses the principle of reweighting ([Calders et al., 2009]. Reweighting of harmed individuals generally matches the intuition of lay users concerning compensatory justice [Velasquez et al., 1990]. To wit, if the weight of an error made on a 28-year old is three times the weight of an error on older clients, the model has to fit its parameters while “paying three times more attention” at previously harmed individuals.
To make the correction of the model fit the lay intuition, FairDream  reweights groups in an ascending way. The weight of an error on a group $S_{k}$ grows with its previous disadvantage. The new sample weight $w_{n}(S_{k})$ of FairDream model #$\#n$ increases with the number of harmed individuals (linearly), and with the difference of fair scores in the baseline model (exponentially). [^8] See Appendix A for more details on FairDream’s correction step. For example, in FairDream’s model #$\#2$, which was selected as the best one, the weight of an error made by the classifier during training on 17-29-year old ($w_{2}(S_{1})$=0.8) became eight times the error on 29-37-year old ($w_{2}(S_{2})$=0.1).

### v3

# Fairness Interventions: A Study in AI Explainability

@@ Abstract
This paper presents a philosophical and experimental study of fairness interventions in AI classification, centered on the explainability of corrective methods. We argue that ensuring fairness requires not only satisfying a target criterion, but also explaining which variables constrain its realization. When corrections are used to mitigate disadvantage transparently, they must remain sensitive to the distribution of true labels. To illustrate this approach, we built FairDream, a fairness package whose mechanism is made transparent for lay users, increasing the model’s weights of errors on disadvantaged groups. While a user may intend to achieve Demographic Parity by the correction method, experiments show that FairDream tends towards Equalized Odds, revealing a conservative bias inherent to the data environment. We clarify the relationship between these fairness criteria, analyze FairDream’s reweighting process, and compare its trade-offs with closely related GridSearch models. Finally, we justify the normative preference for Equalized Odds via an epistemological interpretation of the results, using their proximity with Simpson’s paradox. The paper thus unites normative, epistemological, and empirical explanations of fairness interventions, to ensure transparency for the users.
Keywords: Explainable AI ; Classification ; Fairness ; Equalized Odds ; Demographic Parity; Calibration; GridSearch ; Simpson Paradox; Machine Learning
@@ 1 Introduction
Determining whether an algorithm provides “fair” predictions is a challenging task, although one that is of central practical and ethical importance for society. The difficulties pertain in part to the fact that different notions of fairness exist and compete with each other. On the theoretical side, plethora of statistical metrics exist to quantify the extent to which the predictions of an AI model meet the user’s expectation on fairness, in particular regarding a minority that must not be harmed, [Hellman, 2020] [Mehrabi et al., 2021] [Wan et al., 2023]. Yet, attempts to compare these algorithmic fairness metrics face impossibility theorems, implying that the various fairness criteria proposed cannot be jointly satisfied in all cases (viz. [Kleinberg et al., 2016, Hellman, 2020, Hedden, 2021]).
@@
In this paper, we propose to look at the problem of defining a transparent and explainable AI system for inequality detection and correction. As with other approaches to AI fairness developed in diverse domains, including medicine [Chen et al., 2023], law [Lagioia et al., 2023], or mortgage lending [Lee and Floridi, 2021], we anchor our philosophical analysis to a particular case study, to draw general lessons from the way in which bias can be detected and corrected using AI.
Specifically, we trained an AI model and developed our own fairness package (FairDream) to detect inequalities and then to correct for them. In relation to that, we draw attention to three fairness criteria ([Mayson, 2018, Long, 2021, Hedden, 2021, Barocas et al., 2023]), which can be described qualitatively as follows:
  • Demographic Parity: equalize the overall positive rate across groups.
  • Equalized Odds: equalize true positive rates and false positive rates across groups.
  • Calibration: equalize the ratio of true positives to predicted positives across groups, at identical decision thresholds.
@@
In the rest of this paper, our focus will be on the first two metrics, namely Demographic Parity and Equalized Odds. [^1] Calibration is for us less central to the normative debate concerning the redistribution of resources or true labels. See Appendix C for a discussion of Equalised Odds versus Calibration. The reasons to focus on Demographic Parity and Equalized Odds are twofold. First of all, demands of Demographic Parity are real in various segments of society, in relation to gender, age, or ethnicity. Demographic Parity has thus been interpreted as a legal requirement, e.g. on EU Non-Discrimination Law [Wachter et al., 2020] or for the risk assessment mandated by Pennsylvania [Selbst et al., 2019]. Secondly, our experiments reveal that it is a property of FairDream to fulfill fairness objectives which are conditional on the ground truth (Equalized Odds), even when the user may want to achieve unconditional equality (Demographic Parity). The situation, in that regard, is comparable to the kind of control mechanism found in other intervention systems: to a driver who wants to steer their car, the steering wheel operates in the direction intended by the user, but in ways that take into account other variables (such as speed). Similarly, we explain the way in which the criterion of Equalized Odds constrains such interventions, in particular by comparing it to similar in-processing methods that forego this constraint. Moreover, we propose to explain the discrepancy by analogy to a well-known statistical paradox, the Simpson paradox, to bring to light the right kind of conditional dependency between the predictor variable, the predicted variable, and the true labels.
@@
The paper is structured as follows. In Sections 2 and 3, we first present our training data and the package FairDream. Section 4 introduces the specific property of FairDream of enforcing Equalized Odds, and Section 5 proposes to explain this property by comparison with a closely related approach, the so-called GridSearch method, which can diverge from true labels in order to enforce Demographic Parity. In Section 6, finally, we present some arguments in favor of a metric conditioned on true labels. We respond to the suspicion that the criterion used might be too conservative. We argue that a more revisionary notion of fairness like Demographic Parity too can be enforced, but provided its effects are carefully controlled for. In section 7, we give an explanatory perspective on the correction made by FairDream using an analogy with the Simpson paradox to show the way in which different variables influence each other or not. Finally, in Section 8 we conclude with broader lessons regarding the way in which the explanatory steps taken here make the model trustworthy. We include three appendices: appendix A describes FairDream’s reweighting method, appendix B gives details on the benchmark experiment presented, and appendix C discusses how FairDream performs relative to the metric of Calibration.
@@ 2 Age disparities in the evaluation of income
For our experiment, we thus manipulated the well-known Census dataset, which aggregates data from 48,842 US citizens in 1994.222https://www.kaggle.com/datasets/uciml/adult-census-income. The dataset includes 14 characteristics such as age, work class, years of education, marital status, and geographical origins. A fifteenth column, finally, displays a binary encoding of whether the agent earns less or more than \$50,000 a year (Cf. Figure 1).
@@
Using the Census dataset, we thus built a machine-learning model whose purpose is to minimize the error between its predictions and the actual income of people. Because the practitioner can choose different learners and different metrics of statistical performance, we built a tree-based AI model with the XGBoost library and evaluated it using the standard ROC-AUC measure. [^4] The Receiver Operating Characteristic (ROC) curve plots, for every possible threshold, the tradeoff between its corresponding false positive rate (FPR, x-axis) and true positive rate (TPR, y-axis). Since the ideal binary classifier maximizes TPR and minimizes FPR, the Area Under the ROC Curve (ROC-AUC) is often computed by data scientists as a performance indicator. It reflects along possible thresholds, the ability of the model to rightly select individuals while minimizing positive selection errors. The models of trees ensembles as their evaluation through ROC-AUC are, indeed, broadly used in binary classification on tabular data [Shwartz-Ziv and Armon, 2022]. [^5] An XGBoost classifier, – for eXtreme Gradient Boosted trees,– is an ensemble of trees. Each tree tries to find the best features (splitting populations on their values) to distinguish between individuals earning more or less than \$50,000. Boosting relies on a sequence of weak trees, progressively correcting each others to reach the best predictor. In XGBoost, the computation is made faster through gradient approximation of the errors of the trees, and techniques such as tree-pruning are used to find new trees faster [Chen and Guestrin, 2016]. For stability of the following experiments, we reduced the architecture of the gradient boosted trees to a minimal tuning. By default, we set the model to 1,000 estimators, the maximal depth of each tree being 3 splits on features.
@@ 3 The package FairDream
The main purpose of the package FairDream, originally created in response to an AI fairness contest run by the Monetary Authority of Singapore (MAS), [^6] See for more details the hackathon launched by the MAS under the banner of Veritas initiative. is to give lay users, without any background in data-science or in theories of fairness, an insight into the baseline AI model, and subsequently, a handle on the process of correction. In this section, we present the package FairDream from the user’s perspective. We postpone an explanation of the algorithm’ inner workings and its loss function until the next section.
@@
FairDream takes as input an already trained predictor (the baseline XGBoost we trained) and an indication of the users’ preference for fairness regarding the classificatory variable of interest (i.e. how much the user wants to impose parity regarding the percentage of individuals labeled as $\geq\mathdollar 50k$” by the XGBoost across groups) (see Figure 3). The procedure implemented in FairDream involves two steps, a step of detection and a step of correction.
Figure 3: Menu options in FairDream: fairness objective (top) and disparity factor (bottom)
@@
On the sample of Census used for training, FairDream warns for a ratio of 1 to 5 in the model’s predictions, this ratio being set by the user as sufficient to raise discrimination alerts. For example, for the 17-29 year-old individuals, the portion of the sub-population selected is 12%, against 66% for 29-37 year-old (Cf. Table 1), thereby triggering an alert. The discrimination detection process is fully deterministic and intentionally designed to be transparent and accessible to non-specialist users. Iterative for-loops evaluate selection-rate ratios across fairness metrics for different groups (e.g. gender, age) and trigger an alert whenever these ratios exceed the user-defined discrimination threshold, printed in the intuitive menu of Figure 3.
@@
In our experiment, the stat_score is the indicator of predictive performance chosen by the user, such as ROC-AUC. The fair_score is maximized when the indicators of fairness determined by the user are equalized across groups, such as the overall positive rates for Demographic Parity. We implemented it as a weighted linear sum of the differences of fairness indicators between groups. The best model is then selected as the one with the highest ${\tt trade\_off\_score}=1/3*{\tt stat\_score}+2/3*{\tt fair\_score}$, if for instance the user prefers fairness over accuracy. Note that the selection of the statistical performance metric, fairness criterion, and the balance between these two objectives, is entirely at the discretion of the professional user within FairDream. In our simulations, we reproduce the choices of a user who is not technically specialised but is an expert in banking or insurance, and whose preferences regarding the stat_score or ROC–AUC (see footnote 5) appear legitimate.
Regarding the trade-off between stat_score and fair_score, the weighting coefficients are designed to make it straightforward for the user to prioritize fairness (two-thirds) over statistical performance (one-third). The user may, however, reverse this emphasis or assign equal weights to both. We adopt the fairness-oriented weighting here, since even under this configuration our experiments show that FairDream tends towards Equalized Odds while remaining consistent with the true labels’ environment—analogous to a car’s control system that must account for both steering and speed conditions.
@@
Table 1 shows the predictions of the baseline model at the first step of the procedure, while Table 2 shows the predictions of the model selected as best by FairDream at the second step. The baseline model is trained exclusively on the stat_score objective (ROC–AUC), while the competing models are optimized to balance statistical performance with fairness criteria. Each table highlights the rates of positive predictions (overall positive, true positive, false positive) of the corresponding model’s predictions for two ages groups. The model is evaluated on each age group, first at the global level, then looking at the individuals having the true label $\geq\mathdollar 50,000$’, finally looking at the individuals having the true label $\geq\mathdollar 50,000$’. The labels mentioned in the two tables correspond to the attested events $(Y=1)$ on individuals, i.e. if they really earn over $\mathdollar 50,000$.
At the global level, comparing Table 2 and Table 1 shows that the disparity between age groups in the corrected model is indeed less than 1 to 5, but the correction seems surprisingly small (16 % vs 59 % instead of 12 % vs 66 %), leaving a gap whose ratio is still of 3.7. When comparing by true labels, however, we do see an important correction, since true positive rates and false positive rates come out nearly equal (88% to 89%, 3% to 4%), unlike in the baseline model (68% to 90%, and 2% to 20%). This suggests that the model is not fulfilling Demographic Parity, but a different criterion, known as Equalized Odds. We review the difference between them in the next section.
@@ 4 Demographic Parity vs Equalized Odds
Adopting the taxonomy of [Wachter et al., 2020], one may interpret the objective of diminishing the gap between older and younger clients as an unconditional goal, namely to predict equal proportions of 17-29 and 29-37 year-old as earning over \$50,000, regardless of their true level of income. That would be translated into an objective of “Demographic Parity” [Dwork et al., 2012]: [^10] We notate $Y$ whether an individual actually earns over \$50k, and $\hat{Y}$ the prediction of the model; $A$ and $B$ respectively represent whether an individual is in the 17-29-year old or in the 29-37 year-old category.
@@
This ambiguity calls for explanation, since in the user’s menu options, equalizing the percentage of clients selected as able to repay their loan can be taken to mean that the overall percentages of clients earning over \$50,000 will be predicted as equal across age groups by the FairDream model. Mathematically, this objective corresponds to Demographic Parity. Yet, we see that Demographic Parity is not achieved by the FairDream model. Instead, it actually produces a different result, namely Equalized Odds. Is it an accident? Or is it rather an in-built feature of the model? We show that it is the latter in the next section.
@@ 5.1 An in-processing method
Once we set the goal of equalizing the percentage of clients able to repay their loan across groups, the fairness techniques can equalize them before, during, or after training of the model [Alves et al., 2021]. Although they are implemented in diverse ways, these techniques mainly rely on the following architectural principles:
@@ 5.2 Reweighting for transparent processing
In our attempt to make the method as transparent as possible to lay users, our method uses the principle of reweighting ([Calders et al., 2009]. Reweighting of harmed individuals generally matches the intuition of lay users concerning compensatory justice [Velasquez et al., 1990]. To wit, if the weight of an error made on a 28-year old is three times the weight of an error on older clients, the model has to fit its parameters while “paying three times more attention” at previously harmed individuals. However, in [Calders et al., 2009], the weights of disadvantaged individuals were implemented by duplicating these individuals in the dataset, which was not the most efficient approach, as such duplicated datapoints could be assimilated during training. Our reweighting method instead introduces the weights directly during training, treating them as errors that the classifier learns to avoid with higher priority.
To make the correction of the model fit the lay intuition, FairDream reweights groups in an ascending way. The weight of an error on a group $S_{k}$ grows with its previous disadvantage. The new sample weight $w_{n}(S_{k})$ of FairDream model #n\#n increases with the number of harmed individuals (linearly), and with the difference of fair scores in the baseline model (exponentially). [^11] See Appendix A for more details on FairDream’s correction step. For example, in FairDream’s model #2\#2, which was selected as the best one, the weight of an error made by the classifier during training on 17-29-year old ($w_{2}(S_{1})$=0.8) became eight times the error on 29-37-year old ($w_{2}(S_{2})$=0.1).